«Homework #2 Problem 20.60 MC 6 of Review | Constants Part A Two 2.0-cm-diameter disks are parallel to each other, 2.0 mm apart. They are charged to 10 C What is the electric field strength between the disks? Express your answer using four significant figures. View Available Hint(s) N/C Submit Reguest Answer Part B A proton is placed somewhere between the two disks. What is the magntiude of the electric force acting on that proton due to the electric field between the discs? Express your answer using four significant figures. View Available Hint(s) Submit
Answer #1

a)

electric field between the plates

E = q / ( A * epsilon_not)

E = 10*10^-9 / ( 3.14* 0.01^2* 8.85 10^-12)

E = 3.6*10^6 N/C

=======

b)

F = qE = 1.6*10^-19* 3.6*10^6

F = 5.76*10^-13 N

======

Comment in case any doubt, will reply for sure.. Goodluck